# Prompt-MII: Meta-Learning Instruction Induction for LLMs

### v1

@@ 1 Introduction
There are multiple typical ways to incorporate the given examples: (1) Prompting with instructions, where a natural language task description II is appended to the model prefix, (2) In-context learning (ICL), which directly uses examples in $S_{\text{train}}$ as context during inference, and (3) Supervised fine-tuning (SFT), which performs gradient updates on $S_{\text{train}}$ to condense the information into model parameters. Each method has its advantages. Prompting with instructions is concise and efficient but requires extensive prompt engineering (Sahoo et al., 2024; Schulhoff et al., 2024). ICL achieves highly competitive performance but can be inefficient as the number of examples grows larger (Xiao et al., 2025). SFT is efficient at test time but uses significant compute at training time, requires storage of model weights, and underperforms ICL in many cases (Bertsch et al., 2024).
As a method to bridge the gap between ICL and prompting, there exists instruction induction, which takes training data $S_{\text{train}}$ and generates an instruction II that achieves good performance. Representative methods for instruction induction such as APE (Zhou et al., 2022) and GEPA (Agrawal et al., 2025) typically do so through expensive evolutionary search algorithms at test time that generate multiple candidates for prompts and evaluate them to choose a well-performing prompt. This raises the question: is there a way to perform instruction induction in a way that is both effective and efficient over a wide variety of tasks?
As an answer to this question, we propose Prompt-MII, where we frame instruction induction as a meta-learning problem: instead of individually optimizing II for each individual task, we train an instruction induction policy $\pi_{\theta}$ that can effectively generate instructions in a single pass across diverse task distributions, conditioned only on the in-context examples:
@@
There are two major advantages to this approach. First, it allows $\pi_{\theta}$ to share knowledge about how to construct effective prompts across a wide number of datasets, instead of requiring the re-discovery of this knowledge for each dataset. Second, it has significant efficiency benefits – generating an instruction II for a new dataset simply requires a single forward pass through the language model, instead of a costly optimization process.
@@ 2 Prompt-MII: Meta-learning Instruction Induction
The main challenge in developing a method to generate instructions II from a dataset $S_{\text{test}}$ is learning an effective policy $\pi_{\theta}$ that can generate these instructions in a way that will achieve good test performance. We use reinforcement learning (RL) to train this policy because ground-truth dataset-instruction pairs are often not available for existing public datasets. However, we can evaluate instruction quality through downstream task performance, which serves as a natural reward signal for RL. In this section, we develop our method for meta-learning such a policy, also shown in Figure 2.
@@ 2.1 Training Objective
Let $\mathcal{S}=\{S_{1},S_{2},\ldots,S_{N}\}$ be a collection of datasets that we will use in the meta-learning of $\pi_{\theta}$. For each dataset $S_{i}$, we sample training examples $S_{\text{train}}^{(i)}$ for instruction generation and test examples $S_{\text{test}}^{(i)}$ for reward computation. We define a meta-prompt template $T(S_{\text{train}}^{(i)}))$, which converts the dataset into a prompt to the model, as detailed in subsection $2.$ Then, $\pi_{\theta}$ generates an instruction prompted by this meta-prompt, $I\sim\pi_{\theta}(T(S_{\text{train}}^{(i)}))$.

### v2

@@ 1 Introduction
There are multiple typical ways to incorporate the given examples: (1) Prompting with instructions, where a natural language task description $I$ is appended to the model prefix, (2) In-context learning (ICL), which directly uses examples in $S_{\text{train}}$ as context during inference, and (3) Supervised fine-tuning (SFT), which performs gradient updates on $S_{\text{train}}$ to condense the information into model parameters. Each method has its advantages. Prompting with instructions is concise and efficient but requires extensive prompt engineering (Sahoo et al., 2024; Schulhoff et al., 2024). ICL achieves highly competitive performance but can be inefficient as the number of examples grows larger (Xiao et al., 2025). SFT is efficient at test time but uses significant compute at training time, requires storage of model weights, and underperforms ICL in many cases (Bertsch et al., 2024).
As a method to bridge the gap between ICL and prompting, there exists instruction induction, which takes training data $S_{\text{train}}$ and generates an instruction $I$ that achieves good performance. Representative methods for instruction induction such as APE (Zhou et al., 2022) and GEPA (Agrawal et al., 2025) typically do so through expensive evolutionary search algorithms at test time that generate multiple candidates for prompts and evaluate them to choose a well-performing prompt. This raises the question: is there a way to perform instruction induction in a way that is both effective and efficient over a wide variety of tasks?
As an answer to this question, we propose Prompt-MII, where we frame instruction induction as a meta-learning problem: instead of individually optimizing $I$ for each individual task, we train an instruction induction policy $\pi_{\theta}$ that can effectively generate instructions in a single pass across diverse task distributions, conditioned only on the in-context examples:
@@
There are two major advantages to this approach. First, it allows $\pi_{\theta}$ to share knowledge about how to construct effective prompts across a wide number of datasets, instead of requiring the re-discovery of this knowledge for each dataset. Second, it has significant efficiency benefits – generating an instruction $I$ for a new dataset simply requires a single forward pass through the language model, instead of a costly optimization process.
@@ 2 Prompt-MII: Meta-learning Instruction Induction
The main challenge in developing a method to generate instructions $I$ from a dataset $S_{\text{test}}$ is learning an effective policy $\pi_{\theta}$ that can generate these instructions in a way that will achieve good test performance. We use reinforcement learning (RL) to train this policy because ground-truth dataset-instruction pairs are often not available for existing public datasets. However, we can evaluate instruction quality through downstream task performance, which serves as a natural reward signal for RL. In this section, we develop our method for meta-learning such a policy, also shown in Figure 2.
@@ 2.1 Training Objective
Let $\mathcal{S}=\{S_{1},S_{2},\ldots,S_{N}\}$ be a collection of datasets that we will use in the meta-learning of $\pi_{\theta}$. For each dataset $S_{i}$, we sample training examples $S_{\text{train}}^{(i)}$ for instruction generation and test examples $S_{\text{test}}^{(i)}$ for reward computation. We define a meta-prompt template $T(S_{\text{train}}^{(i)}))$, which converts the dataset into a prompt to the model, as detailed in subsection 2.2. Then, $\pi_{\theta}$ generates an instruction prompted by this meta-prompt, $I\sim\pi_{\theta}(T(S_{\text{train}}^{(i)}))$.